Find the area of the surface generated by revolving the curve y=√25-x², from x = - 2 to x = 3 , about the x-axis.

1
Expert's answer
2022-04-26T11:22:20-0400

"A = 2 \pi \int_a^by \cdot \sqrt{1+(y')^2}" dx

"y = \sqrt{25-x^2}, y' = -\cfrac{x}{\sqrt{25-x^2}}"

"A = 2 \pi \int_{-2}^3 (\sqrt{25-x^2} \cdot\sqrt{1 + \cfrac{x^2}{25-x^2}})dx =\\
=2\pi \int_{-2}^3(\sqrt{25-x^2} \cdot \sqrt{\cfrac{25}{25-x^2}})dx = 2\pi \int_{-2}^3 5 dx = 2\pi \cdot5x|_{-2}^3 = \\
=2\pi (15+10) = 50\pi"